OpenStudy (anonymous):

Let I = [0,1] and let f : I to R be defined by f(x) = x if x is rational and f(x) = 1 - x for x irrational. Show that f is injective on I and that f(f(x))=x and that f is only continuous at the point x = 1/2 Let I = [0,1] and let f : I to R be defined by f(x) = x if x is rational and f(x) = 1 - x for x irrational. Show that f is injective on I and that f(f(x))=x and that f is only continuous at the point x = 1/2 @Mathematics

OpenStudy (alfie):

First, I'd like to say this is an interesting exercise. Let's work together... uhm. First above all some notation. \[\large\begin{array}{l} I = [0,1]\\ f:I \to \mathbb{R} \\ \left\{ \begin{array}{l} f(x) = x\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\; if \;\;\; x \in \mathbb{Q} \\ f(x) = 1 - x \;\;\;\;\;\;\;\;\;\;\;\; if \;\;\;x \in \mathbb{R} \end{array} \right\} \end{array}\] Then we know that a function to be continuous... \[\large \mathop {\lim }\limits_{x \to f({x_0})} f(x) = f({x_0})\] If we think about it: \[\large \begin{array}{l} \mathop {\lim }\limits_{x \to {1^ + }} x \ne 1\\ \mathop {\lim }\limits_{x \to {1^ - }} x \ne 1 \end{array}\] Because in the very moment we do 1+ or 1- we are using an irrational number so of course we have to use the "second part" of the function. But, it is continuos in the 1/2 because: \[\begin{array}{l} \mathop {\lim }\limits_{x \to {{\frac{1}{2}}^ + }} x = 1/2\\ \mathop {\lim }\limits_{x \to {{\frac{1}{2}}^ - }} x = 1/2 \end{array}\] As a matter of fact 1/2 is the only number that subtracted to one gives 1/2.

OpenStudy (alfie):

Now, about injectivity, we have to think about the definition, it means that... \[\large a \ne b \to f(a) \ne f(b)\] We have no kind of problem if they are both rational, because it'd reproduce: \[\large a \ne b\] If just one of them is integer, we find out that : \[\large a \ne 1-a\] and of course: \[\large b \ne 1-b\] What if they are both irrational? Well.. \[\large 1 - a \ne 1 - b \to - a \ne - b \to a \ne b\]

OpenStudy (alfie):

Hope I helped! :) Good luck!

OpenStudy (anonymous):

thnx. but if you approach 1 from the right since I [ 0,1] isn't the right side limit equal to 1?